OpenStudy (anonymous):

√(5 √(5 √(5 √(5)))) solve

mathslover (mathslover):

\(\sqrt{(5\sqrt{(5\sqrt{(5\sqrt{5})))}}}\) Right?

mathslover (mathslover):

\(\begin{array}{}(5\sqrt{5\sqrt{5\sqrt{5}}})^{\cfrac{1}{2}}\\ (5\times 5 \sqrt{5\sqrt{5}})^\cfrac{1}{4}\\ (5\times 5\times 5 \sqrt{5})^\cfrac{1}{8}\\ (5\times 5\times 5\times 5)^\cfrac{1}{16}\\ \end{array}\)

Parth (parthkohli):

\[= \left(5^4\right)^{\large \frac{1}{16}}\]

mathslover (mathslover):

@k.m.imran can you solve it now?

OpenStudy (horotat):

i think its should to be : \[\sqrt{5*\sqrt{5*\sqrt[4]{5^{3}}}}=\sqrt{5*\sqrt[8]{5^{7}}}=\sqrt[16]{5^{15}} = 5^{\frac{ 15 }{ 16 }}\]

OpenStudy (horotat):

hey @mathslover mmm my answer is wrong? :-?

mathslover (mathslover):

Not sure horotat, we need the asker's interaction at this time.

OpenStudy (horotat):

@mathslover at last what is the correct answer?

OpenStudy (anonymous):

its a wrog answer seems

OpenStudy (horotat):

@k.m.imran you mean \[5^{\frac{ 15 }{ 16 }}\] is incorrect?

OpenStudy (anonymous):

for five square root of 5 answer is \[5^{31/32}\]

OpenStudy (horotat):

:-? what is incorect in my answer?

OpenStudy (anonymous):

hey could you explain me how you got that??? @horotat your answer is correct

mathslover (mathslover):

Ok! Horotat, good work dude.

OpenStudy (horotat):

i written my complete answer a little upper :-? you see that?

OpenStudy (anonymous):

HOW DID U GOT

OpenStudy (anonymous):

\[5^{15}\]

mathslover (mathslover):

\(\sqrt{5\sqrt{5 \sqrt{5\sqrt{5}}}}\) Is this the question @k.m.imran ?

OpenStudy (perl):

rewrite using rational exponents starting inside we have (5(5(5(5)^1/2)^1/2)^1/2)^1/2

OpenStudy (anonymous):

yes @mathslover

OpenStudy (anonymous):